# Learning Heterogeneous Mixture of Scene Experts for Large-scale Neural Radiance Fields

### v1

@@ 2 Related Work
Since MoE is typically used to scale up deep learning models, it is often trained on a large number of GPUs. This poses challenges to the efficient implementation of MoE and the designs of large-scale training system. Mesh-TensorFlow [69] provides a multiple node implementation of MoE base on TensorFlow. Deepspeed [70] uses their efficient inference and training system to train MoE at scale and proposes a residual MoE design. Tutel [71] introduces optimized parallelism strategies and CUDA kernels to train large-scale MoE systems. The dispatching operation of Switch-NeRF is based on Tutel kernels while Switch-NeRF designs the full dispatching operation to improve the model accuracy. In Switch-NeRF++, we innovatively merge the dispatching into the hash operation and largely decrease the computation and memory cost of our MoE system.
@@ 3 Switch-NeRF++
Our Switch-NeRF++ learns to encode a large-scale scene into heterogeneous, efficient and scalable Mixture of Hash Experts. Fig. 3 is the overview of the framework in which Fig. 3a and Fig. 3b show the structure of Switch-NeRF and Switch-NeRF++ respectively. Let x denote a 3D point and d denote its direction. x is fed into a gating network $G$. In Switch-NeRF, $G$ is an MLP network to map point features into gate values $G(\textbf{x})$ for point x. In Switch-NeRF++, $G$ is a multi-resolution hash encoding [5]. It looks up a feature encoding $f_{g}$ in the hash encoding by computing a hash value of x. A small MLP uses $f_{g}$ to predict gating values $G(\textbf{x})$ for point x. According to $G(\textbf{x})$, we use Top-$1$ operation to select an expert $E$ and send x into $E$, which is an larger MLP network in Switch-NeRF and a multi-resolution hash encoding with a distinct resolution range in Switch-NeRF++. $E$ computes a feature $f_{e}$ for point x, which is then passed to a small MLP head to predict the point density $\sigma$ and point color c. We add additional appearance embedding AE [29, 2] to capture view-dependent appearance when predicting the color.
@@ 3.1 Trainable Gating Network
The gating network $G$ is pivotal to our method, as it learns the scene decomposition in training. Different previous MoE methods [11, 62] that employs different gating networks in different network layers, our network only leverages one gating network. This design choice stems from the fact that applying more gating networks on massive volume of 3D points in NeRF optimization will largely increase the memory and computation cost. Our experiments also show that more gating networks are not necessary for better decomposition and reconstruction accuracy.
@@
The gating procedure of in Switch-NeRF is:
[…]
The gating procedure of in Switch-NeRF++ is:
[…]
@@
The expert’s final output $\tilde{E}(\textbf{x})$ is computed by multiplying the output of $E_{i}$ with its corresponding gate value:
$\vspace{-1pt}\tilde{E}(\textbf{x})=G(\textbf{x})_{i}E_{i}(\textbf{x}).$ (3)
@@
The use of hash encoding as the gating network in Switch-NeRF++ instead of a pure MLP can provide us with two advantages. The first is that it is much more efficient than a wide MLP head used in previous works [11]. The second is that it is more consistent with our hash experts and the training of gating aligns well with the hash experts. It also produces clearly higher accuracy.
@@ 3.2 Experts for Large Scene Representations
MLP-based NeRF Expert Network. Switch-NeRF employs a collection of $n$ MLP experts$\{E_{i}\}_{i=1}^{n}$, which provide main network parameters to encode scene representation. As illustrated in Fig. 4a, these experts in Switch-NeRF adopt a deep MLP structure with a skip connection, aligning with the core structure design of the vanilla NeRF. The gating network dynamically assigns subsets of 3D points to individual experts. For each 3D point x in an expert $E$, the resulting feature $\tilde{E}(\textbf{x})$ is generated by the expert’s MLP output multiplied by the corresponding gate value. This feature is then passed a unfied prediction head to drive predictions for both the density $\sigma$ and the color c of x.
@@
Unified Head. After generating features of 3D samples from experts, we employ a unified prediction head $H$ to predict the density and color for all the samples. Fig. 4a and Fig. 4b depict the detailed architectures of the unified head in Switch-NeRF and Switch-NeRF++, respectively. The unified head facilitates the sparse framework to encode consistent and unified representations for the different partitions in the entire scene. In the prediction head, a small MLP network is employed to precess the expert output $\tilde{E}(\textbf{x})$ and predict $\sigma$ and color c. The $\tilde{E}(\textbf{x})$ is augmented by concatenating a trainable appearance embedding AE, to model photometric and environmental variations across training images [29].
@@ 3.3 Joint Optimization Losses
Rendering loss. Our large-scale NeRF framework uses the rendering loss for NeRF optimization [1] as primary loss. For a training pixel, we sample $N$ points $\{\textbf{x}_{i}\}_{i=1}^{N}$ along its viewing ray r and predict the density $\sigma_{i}$ and color $\textbf{c}_{i}$ for each point $\textbf{x}_{i}$ by our network. The final color $\hat{C}(\textbf{r})$ is predicted by the volume rendering [1] along the ray r as $\hat{C}(\textbf{r})=\sum_{i=1}^{N}T_{i}(1-\mbox{exp}(-\sigma_{i}\delta_{i}))\textbf{c}_{i},\text{with}~{}T_{i}=\mbox{exp}(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j})$, where $\delta_{i}$ is the distance between neighboring points along r. Let $\mathcal{R}$ be the set of rays in a training batch. The rendering loss $\mathcal{L}_{\textbf{r}}$ is the Mean Squared Error (MSE) between $\hat{C}(\textbf{r})$ and the ground truth $C(\textbf{r})$:
@@
Gating balance loss. Together with $\mathcal{L}_{\textbf{r}}$, we also employ the gating balance loss commonly used in MoE methods [62]. This loss encourages the balanced dispatching of points into different hash experts in order to balance the expert training and utilization. Let $n$ be the expert number and $f_{i}$ be the fraction of points dispatched into an expert $E_{i}$. We have $\sum{f_{i}}=1$. Then, if $\sum{f_{i}^{2}}$ is minimized, the gating is ideally balanced. In order to make $\sum{f_{i}^{2}}$ differentiable for training, as in [62], we replace $f_{i}$ with a soft version $p_{i}$ which is the fraction of gate values dispatched to $E_{i}$. The gating balance loss $\mathcal{L}_{b}$ is defined as $\mathcal{L}_{b}=n\sum{f_{i}g_{i}}.$ $\mathcal{L}_{b}$ is minimized as $1$ if the gating is balanced. Let $\lambda$ be the weight for the gating balance loss, the overall loss $\mathcal{L}_{o}$ is:
@@ 3.4 Unbounded Scene
For MLP-based Switch-NeRF, we use method similar to NeRF++ [28] and Mega-NeRF [2] to divide a large-scale scene into a foreground and a background. In the foreground region, we use cascade sampling similar in NeRF [1]. In the background region, we use the inverse sphere parameterization similar in NeRF++ [28] to handling arbitrary and large depth ranges.
[…]
@@ 3.5 Point dispatching
In the gating operation, after we obtain the expert indexes for the input points, we need to dispatch them into different experts. The MLP-based gating and Hash-based gating are different from each other in the network operation. Therefore, we have different dispatching designs for different gating network to maximum their efficiency.
@@
Capacity factor for training. In MLP-based gating, it is important to efficiently reorganize and send a large number of 3D points into different experts. The dispatching method using Einops-based operations [62] usually struggles with memory overflow due to the massive volume of 3D points. In the training of Switch-NeRF, we adopt the optimized CUDA dispatch operation [71] for training efficiency. We introduce a capacity factor $C_{f}$ that limits the input sample number per NeRF expert in the training, following the setting of previous MoE methods [62]. This ensures uniform tensor dimensions, balanced computation and predictable communication overheads. Given a total batch size $B$, $n$ NeRF experts, the maximum number of sample points per NeRF expert is $B_{e}=\mbox{ceil}(\frac{kBC_{f}}{n})$, where $k$ is the number of chosen experts for each sample.
@@
This uniform dispatch with a capacity factor is depict in Fig. 6. The figure shows that, when $C_{f}=1.0$, overflow points for an expert are dropped and if the expert capacity is not fully used, it will be zero-padded, ensuring uniform tensor shapes. Increasing $C_{f}$ will decreases the dropping ratio but raises memory and computation cost. In the training of Switch-NeRF, we empirically set $C_{f}$ to 1.0 to avoid consuming extra memory. To further mitigate performance loss from lower capacities, we integrate the Batch Prioritized Routing [63] method, which refines training by prioritizing critical samples.
@@
Figure 6: The comparison between the uniform dispatch and full dispatch in Switch-NeRF. In training, the uniform dispatch is utilized for better computation and communication efficiency. The input tensor of each expert has the same shapes. To maintain the shape, it will drop overflow tokens or pad tokens if necessary. In inference, the full dispatch is employed to make sure each point will be processed by an expert, avoiding accuracy decrease. If the uniform dispatch is used in inference, the rendered images will have artifacts.
@@
Full dispatch in inference. While existing MoE methods [62, 12] typically employ the uniform dispatch for both training and testing, this uniform approach may inevitably drops sample points in NeRF inference. In Switch-NeRF’s MLP-based architecture, we find that the uniform dispatch works well in the training but introduces notable accuracy degradation during inference due to dropped sample points. We attribute this to our design of using only one gate network instead of stacked MoE layers and skip connections to maintain the network efficiency. This may make our MoE structure less robust to point dropping in inference. Our solution addressing this problem in inference is to introduce an optimized full dispatch strategy leveraging Pytorch, CUDA, and Tutel. This efficient strategy dispatches all the input points to their corresponding expert with marginal memory overhead, as illustrated in Fig. 6. This strategy yields significant accuracy improvements in final model inference.
3.5.2 Efficient hash dispatching in hash operation gating network.
@@
The MLP gating network use in Switch-NeRF needs extra memory and computation for dispatching because it performs matrix multiplication on features. Different from this, the hash-based gating network fetches features from the hash tables without matrix operations. It first compute a hash index for each 3D points. Then it compute hash table offsets by the hash index and fetch features from hash tables. In our implementation, we find it is very straightforward to we integrate the dispatching into this hash operation. We directly pass the expert indexes into the hash function and use them to compute the hash table offset, which adds only a minimal computation and does not require extra memory to reorganize the inputs. We do not need any explicit rearrange of the input features and do not need to consider capacity factor. Our implementation of hash-based dispatching can further boost the efficiency of our Switch-NeRF++ framework.
@@ 4.1 Datasets and Metrics
Datasets. To thoroughly assess the scalability, accuracy, and efficiency of our method, we conduct experiments on four large-scale datasets, i.e., Mega-NeRF [2], Block-NeRF [3], an adapted version of UrbanBIS [10] and the MatrixCity [73]. The Mega-NeRF dataset comprises 5 scenes (Building, Rubble, Residence, Sci-Art, and Campus), each with $2k$ to $6k$ images of $4k$ to $5k$ resolution, covering areas from 0.15 to 1.3 $km^{2}$. The Block-NeRF dataset includes a scene with $12k$ images at $1k\times 1k$ resolution, covering an $1km$ Urban street. To fully validate the scalability of our method, we adapt a much larger scene (Longhua) from UrbanBIS [10] with $13k$ images at $8k\times 5k$ resolution, covering a $6.5km^{2}$ Urban area. We estimate camera parameters by colmap [37] and downsample the UrbanBIS images by 4. The MatrixCity [73] aerial dataset contains 5 blocks, each with thousands of high-resolution synthetic images. These scenes are very challenging due to the huge image data and the coverage of large-scale regions, suitable for evaluating the scalability of our method.
@@
4.2 Implement Details

### v2

@@ 2 Related Work
Since MoE is typically used to scale up deep learning models, it is often trained on a large number of GPUs. This poses challenges to the efficient implementation of MoE and the design of a large-scale training system. Mesh-TensorFlow [71] provides a multiple-node implementation of MoE based on TensorFlow. Deepspeed [72] uses their efficient inference and training system to train MoE at scale and proposes a residual MoE design. Tutel [73] introduces optimized parallelism strategies and CUDA kernels to train large-scale MoE systems. The dispatching operation of Switch-NeRF is based on Tutel kernels, while Switch-NeRF designs the full dispatching operation to improve the model accuracy. In Switch-NeRF++, we innovatively merge the dispatching into the hash operation and largely decrease the computation and memory cost of our MoE system.
@@ 3 Switch-NeRF++
Our Switch-NeRF++ learns to encode a large-scale scene into heterogeneous, efficient, and scalable Mixture of Hash Experts. Fig. 3 is the overview of the framework in which Fig. 3a and Fig. 3b show the structure of Switch-NeRF and Switch-NeRF++, respectively. Let x denote a 3D point and d denote its direction. x is fed into a gating network $G$. In Switch-NeRF, $G$ is an MLP network to map point features into gate values $G(\textbf{x})$ for point x. In Switch-NeRF++, $G$ is a multi-resolution hash encoding [5]. It looks up a feature encoding $f_{g}$ in the hash encoding by computing a hash value of x. A small MLP uses $f_{g}$ to predict gating values $G(\textbf{x})$ for point x. According to $G(\textbf{x})$, we use Top-$1$ operation to select an expert $E$ and send x into $E$, which is a larger MLP network in Switch-NeRF and a multi-resolution hash encoding with a distinct resolution range in Switch-NeRF++. $E$ computes a feature $f_{e}$ for point x, which is then passed to a small MLP head to predict the point density $\sigma$ and point color c. We add an additional appearance embedding AE [29, 2] to capture view-dependent appearance when predicting the color.
@@ 3.1 Trainable Gating Network
The gating network $G$ is pivotal to our method, as it learns the scene decomposition in training. Different previous MoE methods [11, 64] that employ different gating networks in different network layers, our network only leverages one gating network. This design choice stems from the fact that applying more gating networks on a massive volume of 3D points in NeRF optimization will largely increase the memory and computation cost. Our experiments also show that more gating networks are not necessary for better decomposition and reconstruction accuracy.
@@
The gating procedure of Switch-NeRF is:
$i,G(\textbf{x})_{i}=\mbox{Top-1}(\mbox{Softmax}(\mbox{MLP}(\textbf{x}))),$ (1)
The gating procedure of Switch-NeRF++ is:
$i,G(\textbf{x})_{i}=\mbox{Top-1}(\mbox{Softmax}(\mbox{MLP}(\mbox{Hash}(\textbf{x})))),$ (2)
@@
The expert’s final output $\hat{E}(\textbf{x})$ is computed by multiplying the output of $E_{i}$ with its corresponding gate value:
$\vskip-1.0pt\hat{E}(\textbf{x})=G(\textbf{x})_{i}E_{i}(\textbf{x}).$ (3)
@@
The use of hash encoding as the gating network in Switch-NeRF++ instead of a pure MLP can provide us with two advantages. The first is that it is much more efficient than a wide MLP head used in previous works [11]. The second is that it is more consistent with our hash experts, and the training of gating aligns well with the hash experts. It also produces clearly higher accuracy.
@@ 3.2 Experts for Large Scene Representations
[…]
@@
Unified Head. After generating features of 3D samples from experts, we employ a unified prediction head $H$ to predict the density and color for all the samples. Fig. 4a and Fig. 4b depict the detailed architectures of the unified head in Switch-NeRF and Switch-NeRF++, respectively. The unified head facilitates the sparse framework to encode consistent and unified representations for the different partitions in the entire scene. In the prediction head, a small MLP network is employed to process the expert output $\hat{E}(\textbf{x})$ and predict $\sigma$ and color c. The $\hat{E}(\textbf{x})$ is augmented by concatenating a trainable appearance embedding AE, to model photometric and environmental variations across training images [29].
@@ 3.3 Joint Optimization Losses
Rendering loss. Our large-scale NeRF framework uses the rendering loss for NeRF optimization [1] as the primary loss. For a training pixel, we sample $N$ points $\{\textbf{x}_{i}\}_{i=1}^{N}$ along its viewing ray r and predict the density $\sigma_{i}$ and color $\textbf{c}_{i}$ for each point $\textbf{x}_{i}$ by our network. The final color $\hat{C}(\textbf{r})$ is predicted by the volume rendering [1] along the ray r as $\hat{C}(\textbf{r})=\sum_{i=1}^{N}T_{i}(1-\mbox{exp}(-\sigma_{i}\delta_{i}))\textbf{c}_{i},\text{with}~T_{i}=\mbox{exp}(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j})$, where $\delta_{i}$ is the distance between neighboring points along r. Let $\mathcal{R}$ be the set of rays in a training batch. The rendering loss $\mathcal{L}_{\textbf{r}}$ is the Mean Squared Error (MSE) between $\hat{C}(\textbf{r})$ and the ground truth $C(\textbf{r})$:
@@
Gating balance loss. Together with $\mathcal{L}_{\textbf{r}}$, we also employ the gating balance loss commonly used in MoE methods [64]. This loss encourages the balanced dispatching of points into different hash experts in order to balance the expert training and utilization. Let $n$ be the expert number and $f_{i}$ be the fraction of points dispatched into an expert $E_{i}$. We have $\sum{f_{i}}=1$. Then, if $\sum{f_{i}^{2}}$ is minimized, the gating is ideally balanced. In order to make $\sum{f_{i}^{2}}$ differentiable for training, as in [64], we replace $f_{i}$ with a soft version $p_{i}$ which is the fraction of gate values dispatched to $E_{i}$. The gating balance loss $\mathcal{L}_{b}$ is defined as $\mathcal{L}_{b}=n\sum{f_{i}g_{i}}.$ $\mathcal{L}_{b}$ is minimized as $1$ if the gating is balanced. Let $\lambda$ be the weight for the gating balance loss; the overall loss $\mathcal{L}_{o}$ is:
@@ 3.4 Unbounded Scene
For MLP-based Switch-NeRF, we use a method similar to NeRF++ [28] and Mega-NeRF [2] to divide a large-scale scene into a foreground and a background. In the foreground region, we use cascade sampling similar to NeRF [1]. In the background region, we use the inverse sphere parameterization similar to NeRF++ [28] to handle arbitrary and large depth ranges.
For hash-based Switch-NeRF++, the hash encodings are defined on a grid in a bounded 3D space. In the foreground region, we use the normal hash grids. The background uses a contraction method in Mip-NeRF 360 [74] to map the unbounded space into a bounded space. In the contraction, we normalize the scene into a unit sphere with the scene bound, and then apply the contraction on a 3D point x. If $\|\textbf{x}\|\leq 1$, we do not change its value. If $\|\textbf{x}\|>1$, we transform $\|\textbf{x}\|$ by $\left(2-\frac{1}{\|\textbf{x}\|}\right)\left(\frac{\textbf{x}}{\|\textbf{x}\|}\right)$. This contraction function maps the background 3D points into a bounded scene without affecting the 3D points in the foreground. The hash grids for the background are defined in the contracted space.
@@ 3.5 Point dispatching
In the gating operation, after we obtain the expert indexes for the input points, we need to dispatch them to different experts. The MLP-based gating and the hash-based gating are different from each other in the network operation. Therefore, we have different dispatching designs for different gating network to maximize their efficiency.
@@
Capacity factor for training. In MLP-based gating, it is important to efficiently reorganize and send a large number of 3D points to different experts. The dispatching method using Einops-based operations [64] usually struggles with memory overflow due to the massive volume of 3D points. In the training of Switch-NeRF, we adopt the optimized CUDA dispatch operation [73] for training efficiency. We introduce a capacity factor $C_{f}$ that limits the input sample number per NeRF expert in the training, following the setting of previous MoE methods [64]. This ensures uniform tensor dimensions, balanced computation, and predictable communication overheads. Given a total batch size $B$, $n$ NeRF experts, the maximum number of sample points per NeRF expert is $B_{e}=\mbox{ceil}(\frac{kBC_{f}}{n})$, where $k$ is the number of chosen experts for each sample.
@@
This uniform dispatch with a capacity factor is depicted in Fig. 6. The figure shows that, when $C_{f}=1.0$, overflow points for an expert are dropped, and if the expert capacity is not fully used, it will be zero-padded, ensuring uniform tensor shapes. Increasing $C_{f}$ will decrease the dropping ratio but raise memory and computation cost. In the training of Switch-NeRF, we empirically set $C_{f}$ to 1.0 to avoid consuming extra memory. To further mitigate performance loss from lower capacities, we integrate the Batch Prioritized Routing [65] method, which refines training by prioritizing critical samples.
@@
Figure 6: The comparison between the uniform dispatch and full dispatch in Switch-NeRF. In training, the uniform dispatch is utilized for better computation and communication efficiency. The input tensor of each expert has the same shape. To maintain the shape, it will drop overflow tokens or pad tokens if necessary. In inference, the full dispatch is employed to make sure each point will be processed by an expert, avoiding accuracy decrease. If the uniform dispatch is used in inference, the rendered images will have artifacts.
@@
Full dispatch in inference. While existing MoE methods [64, 12] typically employ the uniform dispatch for both training and testing, this uniform approach may inevitably drop sample points in NeRF inference. In Switch-NeRF’s MLP-based architecture, we find that the uniform dispatch works well in the training but introduces notable accuracy degradation during inference due to dropped sample points. We attribute this to our design of using only one gate network instead of stacked MoE layers and skip connections to maintain the network efficiency. This may make our MoE structure less robust to point dropping in inference. Our solution addressing this problem in inference is to introduce an optimized full dispatch strategy leveraging Pytorch, CUDA, and Tutel. This efficient strategy dispatches all the input points to their corresponding expert with marginal memory overhead, as illustrated in Fig. 6. This strategy yields significant accuracy improvements in final model inference. The full dispatch operation is also explored in recent large language models (LLMs) based on MoE. Deepseek-V3 [75] describes a similar strategy, where the model maintains good load balance and avoids token dropping throughout both training and inference. This approach aligns closely with our motivation for using full dispatch at inference time, especially in the context of 3D rendering.
3.5.2 Efficient hash dispatching in hash gating network.
@@
The MLP gating network used in Switch-NeRF needs extra memory and computation for dispatching because it performs matrix multiplication on features. Different from this, the hash-based gating network fetches features from the hash tables without matrix operations. It first computes a hash index for each 3D point. Then it computes hash table offsets by the hash index and fetches features from the hash tables. In our implementation, we find it is very straightforward to integrate the dispatching into this hash operation. We directly pass the expert indexes into the hash function and use them to compute the hash table offset, which adds only minimal computation and does not require extra memory to reorganize the inputs. We do not need any explicit rearrangement of the input features and do not need to consider the capacity factor. Our implementation of hash-based dispatching can further boost the efficiency of our Switch-NeRF++ framework.
3.6 Method conclusion.
Figure 7: Block diagram illustrating the integrated architectural innovations in Switch-NeRF++. With these innovations, our method exhibits significant efficiency gains, with an 8x acceleration in training and a 16x acceleration in rendering compared to Switch-NeRF.
In conclusion, Switch-NeRF++ integrates a set of architectural innovations that collectively improve both accuracy and efficiency over the original Switch-NeRF. As shown in Fig. 7, the efficient hash gating network combined with full dispatching ensures both the speed and accuracy of expert selection. The heterogeneous hash experts provide accurate and compact scene representations. The innovations in gating, dispatching, and experts play a critical role in speeding up the training and inference. The unified prediction head (MLP) ensures consistency across different experts. Together with these innovations, the Switch-NeRF is upgraded into Switch-NeRF++ with substantial improvements in efficiency and accuracy.
@@ 4.1 Datasets and Metrics
Datasets. To thoroughly assess the scalability, accuracy, and efficiency of our method, we conduct experiments on four large-scale datasets, i.e., Mega-NeRF [2], Block-NeRF [3], an adapted version of UrbanBIS [10], and the MatrixCity [76]. The Mega-NeRF dataset comprises 5 scenes (Building, Rubble, Residence, Sci-Art, and Campus), each with $2k$ to $6k$ images of $4k$ to $5k$ resolution, covering areas from 0.15 to 1.3 $km^{2}$. The Block-NeRF dataset includes a scene with $12k$ images at $1k\times 1k$ resolution, covering a $1km$ Urban street. To fully validate the scalability of our method, we adapt a much larger scene (Longhua) from UrbanBIS [10] with $13k$ images at $8k\times 5k$ resolution, covering a $6.5km^{2}$ Urban area. We estimate camera parameters by colmap [37] and downsample the UrbanBIS images by 4. The MatrixCity [76] aerial dataset contains 5 blocks, each with thousands of high-resolution synthetic images. These scenes are very challenging due to the huge image data and the coverage of large-scale regions, suitable for evaluating the scalability of our method.
@@
4.2 Implementation Details
Algorithm 1 The forward and loss computation of Switch-NeRF++.
0:Inputs: 3D points $\mathbf{x}$, view directions $\mathbf{d}$, appearance embeddings $\mathbf{a}$
1:// === Part I: Network Configuration ===
2:// — Hash Gating Network —
3: Define base_resolution: $16$
4: Define desired_resolution: $2048$
5: Define hashmap_size: $2^{19}$
[…]
7:// — Hash Experts —
8: Define number of hash experts $N=8$
9: Define a set of base_resolution $\{r\}$, logarithmically spaced in $[16,512]$ for $N=8$ hash experts
10: Define a set of desired_resolution $\{\tilde{r}\}$, logarithmically spaced in $[2048,16384]$ for $N=8$ hash experts
11:for $i=1$ to NN do
12:  $\text{HashExpert}_{i}=\text{HashEncoding}({\{r\}_{i}}^{3},{\{\tilde{r}\}_{i}}^{3})$
13:end for
14: The resolution intervals are logarithmically sampled to cover the scale distribution of urban scenes.
15:// === Part II: Forward Pass ===
16:// — Gating Network —
17: Encode position: $\mathbf{f}_{g}\leftarrow\text{HashGate}(\mathbf{x})$
18: Compute gate values: $\mathbf{G}=\text{Softmax}(\mathbf{f}_{g})$
19: Select top-1 hash expert: $i=\arg\max(\mathbf{G})$
20:// — Hash Expert Processing —
21: Hash Expert encoding: $\mathbf{E}_{i}(\mathbf{x})\leftarrow\text{HashExpert}_{i}(\mathbf{x})$
22: Apply gate value: $\hat{\mathbf{E}}_{i}(\mathbf{x})=\mathbf{G}_{i}\cdot\mathbf{E}_{i}(\mathbf{x})$
23:// — Prediction —
24: Direction encoding: $\mathbf{SH}_{d}=\text{SphericalHarmonics}(\mathbf{d})$
25: Concatenate: $\mathbf{f}_{\text{final}}=[\hat{\mathbf{E}}_{i}(\mathbf{x}),\mathbf{SH}_{d},\mathbf{a}]$
26: Predict density: $\sigma=\text{MLP}_{\text{density}}(\hat{\mathbf{E}}_{i}(\mathbf{x}))$
27: Predict color: $\mathbf{c}=\text{MLP}_{\text{color}}(\hat{\mathbf{E}}_{i}(\mathbf{x}))$
28:// Rendering
[…]
30: Rendered Color: $\hat{C}(\textbf{r})=\sum_{i=1}^{N}T_{i}(1-\mbox{exp}(-\sigma_{i}\delta_{i}))\textbf{c}_{i}$
31:// Losses
32: Ground Truth Color: $C(\textbf{r})$
33: Rendering Loss: $\mathcal{L}_{\textbf{r}}=\sum_{\textbf{r}\in\mathcal{R}}\left\|\hat{C}(\textbf{r})-C(\textbf{r})\right\|_{2}^{2}$
34: Gate Balance Loss: $\mathcal{L}_{b}=n\sum{f_{i}g_{i}}.$
35: Balance Loss Weight: $\lambda=5\times 10^{-4}$
36: Overall loss: $\mathcal{L}_{o}=\mathcal{L}_{\textbf{r}}+\lambda\mathcal{L}_{b}$